Presentation is loading. Please wait.

Presentation is loading. Please wait.

SPS LS1 Planning meeting (N017) 17/07/13 David Mcfarlane

Similar presentations

Presentation on theme: "SPS LS1 Planning meeting (N017) 17/07/13 David Mcfarlane"— Presentation transcript:

1 SPS LS1 Planning meeting (N017) 17/07/13 David Mcfarlane

2 Topics Lifts and Monte-charges. TT10 update. Removal of equipment in LSS1+ Works around the CCC (Surface) Positionnement grues BA3. Planning. False floors open in the BAs E.C.Rs. Access rights and safety equipment Important Information. A.O.B

3 Lift Maintenance July/August SPS.Technical-Coordination@Cern.ch3

4 TT10 The first test support was installed last week. SPS.Technical-Coordination@Cern.ch4

5 New storage location for Vacuum chambers SPS.Technical-Coordination@Cern.ch5

6 Removal of equipment in LSS1+ Work has started to remove the equipment in LSS1+ in preparation for the irradiated cable campaign that will start in October. The most radioactive equipment will be removed first (Dumps etc) When they are moved round the tunnel to ECX5 all of the zones that transport will pass through MUST be clear of equipment AND personnel. (BA1-, BA6 and BA5+) When the exact times of transport are known, everyone will be informed and someone from coordination will pass in front of transport to ensure the areas are clear. All equipment owners that are transporting equipment must make and EDH request for transport. SPS.Technical-Coordination@Cern.ch6

7 Works around the CCC Installation of new Transformers Creation of the new cable trench from the CCC to BA3 Works inside Building 874 (this includes works on the outside of the building. Access required around CCC for HVAC replacement works Installation of PIPES trench between CCC and 774 Installation of CABLE trench between CCC and 774 Installation of WASTE WATER trench between CCC and 774 (near Jura) Waste water pump and pipe work (Seleve side of CCC) Changing of SPS ring-main valves Fibre Optics Works (Jura side of CCC) SPS.Technical-Coordination@Cern.ch7

8 8

9 Planning for the next few months SPS.Technical-Coordination@Cern.ch9

10 Positionnement grues BA3 Salut David, Nous commencerons donc les transfos au BA3 la semaine prochaine. Nous essayerons de perturber le moins possible les alentours de la CCC mais il faudra quand même bloquer temporairement la route Adams Au niveau des accès vers les transfos (équipement de la grue) plus les Allés et venues des remorques (8 transfos -> 16 passages..).. Je te tiendrai informé en fonction de lavancement Bonne journée. Serge SPS.Technical-Coordination@Cern.ch10


12 Dépose de tous les anciens transformateurs Possible avec la Grue 55 t… ~10 m

13 Dépose de tous les anciens transformateurs Possible avec la Grue 55 t…(voir test a blanc)

14 Mise en place transformateurs Dipôles Grue 55 t




18 Mise en place transformateurs Quadripôles Grue 220 t




22 EN UndergroundSur BA1 UndergroundSur BA2 UndergroundSur BA3 UndergroundSur BA4 UndergroundSur BA5 UndergroundSur BA6 TI8TI2 TRANSFER TUNNELS UndergroundSur BA7 UndergroundSur BA80/TCC2 2013 2014 2015 Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. 2013 2014 2015 Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. May Jun. Jul. Aug. Sep. Oct. Nov. Dec. Jan. Feb. Mar. Apr. General maintenance Cold check out & Start-up DSO tests/HV tests EPC Commissioning tests Vacuum Ions to NA61 General Maintenance General Maintenance General Maintenance General maintenance General Maintenance UA9 works RF upgrade works EPC works Civil Engineering RF upgrade works RF Upgrade works Survey Horizontal Vertical smoothing Cables Irradiated Cables Cables Re-route Cabling Cables EL Studies Prep work in LSS1 Re-install equipment removed for cable works Survey SPS LS1 schedule Beam on Cool down/RP survey and Magnet tests Protons to the North Area Protons to LHC Cool down NO ACCESS Cool down Limited ACCESS Magnets Beam off : 16/02/13 Beam on: 13/10/14 Open Access : 27/03/13 Close Access : 27/06/14 Fibres CV works CV studies CV Controls upgrade Vacuum TT10 Consolidation 18 KV Septa works Kicker Septa works Kicker TT20 pipework Magnets 18 KV Civil Engineering Magnets Ring-main Vacuum Crab Cavities EPC works Cooling tower 863 Ring-main Cables Fibres 18 KV Civil Engineering in TT10 Fibres CV maintenance Cables Kicker Cables RAL Septa replacement

23 False floors lifted in Surface BAs BAsCabling pulling startCable pulling end BA1 (TS1- works) BA1 (Irradiated cable works) BA1 (Fibre Optics) 29/04/2013 08/10/2013 08/04/2013 02/08/2013 14/02/2014 24/04/2013 BA206/01/201411/03/2014 BA3 BA3 (Fibre Optics) 27/03/2013 22/04/2014 31/05/2013 06/05/2014 BA427/05/201312/07/2013 BA5 BA5 (Fibre Optics) 27/03/2013 30/09/2013 24/05/3013 11/10/2013 BA6 BA6 (Fibre Optics) 30/09/2013 13/5/2013 18/10/2013 24/05/2013 BA716/10/201307/11/2013 SPS.Technical-Coordination@Cern.ch23 Below is a table showing the dates that EL will be pulling cables at each point of the SPS. During these periods large sections of the false floor in these surface buildings will be lifted. Any works in the BAs that involve lifting of the false floors MUST be reported to the Coordination team If you require access to racks or other equipment in the surface buildings during these periods you must check with planning first.

24 Engineering Change Requests Templates and Guidelines for Engineering Change Requests: – EDMS: 12718801271880 SPS ECR template (in DFS): – pplications/CERN/Workgroup%20Templates/SPS/ECR-Template- SPS.dotx pplications/CERN/Workgroup%20Templates/SPS/ECR-Template- SPS.dotx SPS.Technical-Coordination@Cern.ch24

25 Current ECRs for LS1 1.Removal of AEWB Pick-up from position 31932 to 31502 (EDMS No. 1245361) RELEASED 2.Removal of a BPWA Pick-up from position 31931 to 31101 (EDMS No. 1245393) RELEASED 3.Installation of a new Wideband Kickers system in position 32173 (EDMS No. 1245400) RELEASED 4.Installation of a new HBW Damper Pick-up in place of BPWA 31931 & AEWB 31932 (EDMS No. 1245401) RELEASED 5.Replacement of 4 old BPA and 2 BPD by 6 BPCE in LSS1 & LSS2 (EDMS No. 1246559) RELEASED 6.Reinstallation of pump and gauges (Pirani + Penning) in 11993 (vacuum sector 150) (EDMS No. 1249305) RELEASED 7.Modification LSS1 sectorization (new VVSF + displacement VVSB 11772) (EDMS No. 1249307 v.2) RELEASED 8.Installation of 3 BCT in LSS5 for North Area protection during Ions operations in SPS. (EDMS No. 1075945) UNDER IMPLEMENTATION SPS.Technical-Coordination@Cern.ch25

26 Current ECRs for LS1 9.Transfer of the Fast Beam Current Transformers (BCTFR.31451) for the SPS circulating beam from LSS3 to LSS5. (EDMS No. 1221341) RELEASED 10.SPS Beam Position Monitoring Prototyping Setup in LSS4. (EDMS No. 1099630) RELEASED 11.Installation de 4 moniteurs de position (2 BPH, 2 BPV), dédiés au système Damper. (EDMS No. 1251693) RELEASED 12.Installation d'un nouveau moniteur de position de type BPCR en LSS2 au 22105. (EDMS No. 1251695) RELEASED 13.Displacement of Door PPG2145 separating the LHC and the SPS in the TI2 tunnel. (EDMS No. 1253832) RELEASED 14.Interlocking of Injection Tunnel upstream areas with LHC powering. (EDMS No. 1257949) RELEASED 15.Move of BPH.31998 to BPH.42173. (EDMS No. 1272718) RELEASED 16.Optical Fibre Racks. (EDMS No. 1273766) RELEASED SPS.Technical-Coordination@Cern.ch26

27 Current ECRs for LS1 17.Removal of old fixed TMP vacuum groups in 11655-11691. (EDMS No. 1273232) RELEASED 18.BYBPM racks under the SPS dipole magnets as part of the Upgrade of Beam Position Monitors electronics (MOPOS). (EDMS No. 1282651) UNDER IMPLEMENTATION 19.Installation de Coffrets d'Alimentation en LSS1 et TT10. (EDMS No. 1282860) RELEASED 20.Installation d'un rack de jonction RX11695=TS1 au BA1, afin de limiter les longueurs des câbles à changer lors d'un futur remplacement. (EDMS No. 1288607) RELEASED 21.Installation de nouveaux chemins de câbles en LSS1+ et LSS2- du SPS. (EDMS No. 1291889 v.2) IN WORK SPS.Technical-Coordination@Cern.ch27

28 Access rights and safety equipment Everyone who enters the SPS MUST have the following: – Completed the correct courses. Safety course levels 1,2 and 3. Habilitation Electrique. Radiological Protection. Self Rescue Mask Training. – The correct access authorisation (via EDH). – A valid approved IMPACT. (with their name included on this IMPACT) – They must know the number of this IMPACT. – Safety shoes. – Biocell. – Electronic dosimetre. – Passive dosimetre. – Helmet plus Light. Please note that the medical service no longer contact you when your medical is about to expire!!!!!!! – You will receive an e-mail 2 months before informing you that some of your access rights will soon expire. – It is up to you to contact the medical service to arrange a medical. SPS.Technical-Coordination@Cern.ch28

29 Important information Useful Contact Numbers – SPS Coordination Team David MCFARLANE- 164247 Franck BAIS – 165467 James RIDEWOOD - 160590 – Control Room SPS CCC – 77500 TI CCC – 72201 CSA - 62511 – RP Christophe Tromel – 163199 Angelito Herve – 163168 Nadine Conan (BA1) – 160641 Kurt Weiss (BA1) -160759 – Coordination Safety Emmanuel Paulat – 163870 John Etheridge - 164647 SPS.Technical-Coordination@Cern.ch29

30 There are currently 47 IMPACTs that are LATE. If you have completed your work then you need to close the related IMPACT. I will be absent from CERN on Friday the 19 th July and Monday the 22 nd July. You can contact Franck or James on matter concerning the SPS. A.O.B. SPS.Technical-Coordination@Cern.ch30

31 SPS coordination website \\\J\jocksoft\Public\LS1 Next meeting 31/07/2013 892-1-D20 SPS.Technical-Coordination@Cern.ch31

Download ppt "SPS LS1 Planning meeting (N017) 17/07/13 David Mcfarlane"

Similar presentations

Ads by Google